OpenStudy (anonymous):

A small software company has three customer service representatives. After a week of observation, the supervisor of the customer service department determined that there is an 85% probability that a customer service representative will be on the phone with a customer at any given time. What is the probability of all three representatives being on the phone at the same time? Round to the nearest percent if necessary.

OpenStudy (anonymous):

P(3 on phone) = 0.85^3 =?

OpenStudy (anonymous):

can u please calculate

OpenStudy (anonymous):

\[0.85^{3} =0.6141\]

OpenStudy (anonymous):

changing that percentage

OpenStudy (anonymous):

\[0.6141 \times 100 = 61.41 \]

OpenStudy (anonymous):

so 61.41 %

OpenStudy (anonymous):

Thank you!!